# Characteristic Learning for Provable One Step Generation

### v3

@@ 3.3. Analysis for Velocity Matching
As shown in Proposition 2.3, the $2$-Wasserstein error is bounded by the averaged $L^{2}$-risk of the velocity estimator. The majority of existing literature on theoretical analysis of diffusion and flow-based generative models commonly assumes that the $L^{2}$-risk of score or velocity matching is sufficiently small (Lee et al., 2022, 2023; Chen et al., 2023d, c; Benton et al., 2024a, b; Gao and Zhu, 2024). However, this line of research lacks the ability to quantitatively characterize the convergence rate of velocity matching with respect to the number of samples. Additionally, no prior theoretical guidance for the selection of neural networks is provided in this literature. To the best of our knowledge, only a limited number of studies have specifically focused on investigating the convergence rates of score matching (Oko et al., 2023; Chen et al., 2023b; Han et al., 2024) and velocity matching (Chang et al., 2024; Gao et al., 2024; Jiao et al., 2024).
@@
=${\mathscr{B}}=\left\{b\in N(L,S):\begin{aligned} &\|b(t,x)\|_{\infty}\leq B_{\operatorname{vel}}\log^{1/2}n,\\ &\|\partial_{t}b(t,x)\|_{2}\leq 3D\kappa(T)\log^{1/2}n,\\ &\|\nabla b(t,x)\|_{\operatorname{op}}\leq 3G,\leavevmode\nobreak\ (t,x)\in[0,T]\times{\mathbb{R}}^{d}\end{aligned}\right\},$
where the depth and the width of the neural network are given, respectively, as $L=C$ and $S=C\kappa^{\frac{2(d+1)}{d+3}}(T)n^{\frac{d+1}{d+3}}$. Then the following inequality holds
@@
The rate of velocity matching in Theorem 3.9 is consistent with the minimax optimal rate ${\mathcal{O}}(n^{-\frac{2}{d+3}})$ in nonparametric regression (Stone, 1982; Yang and Barron, 1999; Gyorfi et al., 2002; Tsybakov, 2009) given that the target function is Lipschitz continuous. Moreover, our theoretical findings align with convergence rates of nonparametric regression using deep neural networks (Bauer and Kohler, 2019; Nakada and Imaizumi, 2020; Schmidt-Hieber, 2020; Kohler and Langer, 2021; Farrell et al., 2021; Kohler et al., 2022; Jiao et al., 2023a). It is noteworthy that our results improve upon the rate ${\mathcal{O}}(n^{-\frac{2}{d+5}})$ derived by Chen et al. (2023b); Chang et al. (2024).
In Theorem 3.9, the hypothesis class ${\mathscr{B}}$ is set as a deep neural network class with Lipschitz constraints. The assumption of uniformly Lipschitz continuity of velocity estimator plays a crucial role in controlling the discretization error induced by Euler method. For further details, refer to Theorem 3.11. This assumption is standard in the theoretical analysis for flow-based or diffusion models, as considered by Kwon et al. (2022, Assumption A2) and Chen et al. (2023c, Assumption 3). In practical applications, various techniques for restricting the Lipschitz constant of deep neural networks have been reviewed in Section 2.2. In the theoretical perspective, the approximation properties of deep neural network with Lipschitz constraint has been studied by Chen et al. (2022); Huang et al. (2022); Jiao et al. (2023b); Ding et al. (2024). In this work, an approximation error bound for deep neural networks with Lipschitz constraint is established in Section H.
With the aid of Proposition 2.3 and Theorem 3.9, we obtain a 2-Wasserstein error of the estimated flow (2.9) as follows.
@@ 3.4. Analysis for Euler Sampling
Under the same conditions as Theorem 3.9. Let the number of time steps for Euler method $K$ be a positive integer. Then the following inequality holds
@@
The error bound of the Euler flow (2.10), as derived in Theorem 3.11, can be roughly divided into two main components. The first term arises from velocity matching, aligning with the error bound presented in Corollary 3.10. The second term corresponds to the discretization error introduced by Euler method. Moreover, as the number of time steps $K$ increases, the error bound in Theorem 3.11 converges to that in Corollary 3.10.
@@
Under the same conditions as Theorem 3.11. Suppose the number of time steps for Euler method $K$ satisfies $K\geq Cn^{\frac{1}{d+3}}$. Then the following inequality holds
@@
Corollary 3.12 presents a bound for the total error of flow-based generative models. The first term in the error bound corresponds to the 2-Wasserstein error of the Euler flow, as derived in Theorem 3.11. The second term captures the convergence of the interpolant distribution $\mu_{T}$ to the target distribution $\mu_{1}$. It is worth noting that as the stopping time $T$ approaches one, the first term tends to infinity, while the second term simultaneously decreases. This trade-off within the error bound highlights the importance of carefully selecting the stopping time $T$ and provides practical guidance for determining it in real-world applications.
@@ 3.5. Analysis for Characteristic Generator
[…]
where the depth and the width of the neural network are given, respectively, as $L=C$ and $S=Cm^{\frac{d+2}{d+4}}$. Then it follows that
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{Z}}}\big{[}{\mathcal{D}}(\widehat{g})\big{]}\leq C\kappa^{2}(T)n^{-\frac{2}{d+3}}\log^{2}n+Cm^{-\frac{2}{d+4}}\log^{2}m+C\Big{(}\frac{\log m}{K}+\frac{\log n}{K^{2}}\Big{)},$
@@
(3.3) $K\geq\max\big{\{}Cn^{\frac{1}{d+3}},C\kappa^{-2}(T)n^{-\frac{2}{d+3}}\big{\}}\quad\text{and}\quad m\geq C\kappa^{-(d+4)}(T)n^{\frac{d+4}{d+3}},$
@@ 3.6. Applications
In this section, the theoretical analysis is applied to two types of flow-based method in Table 2: linear interpolant and Föllmer flow. The convergence rates of them are shown in Corollaries 3.14 and 3.15, respectively.
@@
Under the same conditions as Theorem 3.13. Set the stopping time $T$ as
@@
Further, suppose the number of time step $K$ and number of samples $m$ satisfies (3.3). Then the following inequality holds
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{Z}}}\big{[}{\mathcal{D}}(\widehat{g})\big{]}\leq Cn^{-\frac{2}{3(d+3)}}\log n,$
where the constant $C$ is independent of $n$.
@@
Under the same conditions as Theorem 3.13. Set the stopping time $T$ as
@@
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}W_{2}^{2}\Big{(}(\widehat{E}_{0,K}^{\tau})_{\sharp}\mu_{0},\mu_{1}\Big{)}\Big{]}\leq CCn^{-\frac{2}{5(d+3)}}\log^{\frac{3}{5}}n.$
@@
Further, suppose the number of time step $K$ and number of samples $m$ satisfies (3.3), respectively. Then the following inequality holds
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{Z}}}\big{[}{\mathcal{D}}(\widehat{g})\big{]}\leq Cn^{-\frac{2}{5(d+3)}}\log^{\frac{3}{5}}n,$
where the constant $C$ is independent of $n$.
@@ 4.1.3. Characteristic fitting
8:     ^$\widehat{{\mathcal{R}}}_{\operatorname{loc}}(D_{{\mathcal{S}},\phi})\leftarrow\|D_{{\mathcal{S}},\phi}(t,t,X_{t})-X_{0}\|_{2}^{2}$.
@@
11:     ^$\widehat{{\mathcal{R}}}_{\operatorname{glo}}(D_{{\mathcal{S}},\phi})\leftarrow\|g_{s,T}^{\operatorname{off}}\circ g_{u,s}\circ g_{t,u}^{\operatorname{int}}(X_{t})-g_{s,T}^{\operatorname{off}}\circ g_{t,s}^{\operatorname{off}}(X_{t})\|_{2}^{2}$.
@@
Kim et al. (2024) proposed a similar method, but our approach differs from the CTM in two significant aspects. First, the integral scheme $g^{\operatorname{int}}$ employed by CTM is Euler-based. While Euler method coincides with the first-order exponential integrator for VE-ODE (Song et al., 2021c), its numerical stability cannot be guaranteed for general probability flow ODEs. In contrast, the exponential integrator used in our method may ease potential training instability as it fully exploits the semi-linearity of the ODE system. Secondly, CTM relies on GAN training in their approach, borrowing a pre-trained discriminator and treating $g$ as the generator. This reliance on GAN training may cause potential training instability and require extra training of the discriminator. Our method, however, does not require this additional GAN training burden.

### v5

@@ 3.3. Analysis for Velocity Matching
The majority of existing literature on theoretical analysis of diffusion and flow-based generative models commonly assumes that the $L^{2}$-risk of score or velocity matching is sufficiently small (Lee et al., 2022, 2023; Chen et al., 2023d, c; Benton et al., 2024a, b; Gao and Zhu, 2024; Beyler and Bach, 2025). However, this line of research lacks the ability to quantitatively characterize the convergence rate of velocity matching with respect to the number of samples. Additionally, no prior theoretical guidance for the selection of neural networks is provided in this literature. To the best of our knowledge, only a limited number of studies have specifically focused on investigating the convergence rates of score matching (Oko et al., 2023; Chen et al., 2023b; Han et al., 2024) and velocity matching (Chang et al., 2024; Gao et al., 2024; Jiao et al., 2024).
@@
${\mathscr{B}}=\left\{b\in N(L,S):\begin{aligned} &\|b(t,x)\|_{\infty}\leq B_{\operatorname{vel}}\log^{1/2}n,\\ &\|\partial_{t}b(t,x)\|_{2}\leq 3D\kappa(T)\log^{1/2}n,\\ &\|\nabla b(t,x)\|_{\operatorname{op}}\leq 3G,\penalty 10000\ (t,x)\in[0,T]\times{\mathbb{R}}^{d}\end{aligned}\right\},$
where the depth and the number of non-zero weights of the neural network are given, respectively, as $L=C$ and $S=Cn^{\frac{d+1}{d+3}}$. Then the following inequality holds
@@
The rate of velocity matching in Theorem 3.10 is consistent with the minimax optimal rate ${\mathcal{O}}(n^{-\frac{2}{d+3}})$ in nonparametric regression (Stone, 1982; Yang and Barron, 1999; Gyorfi et al., 2002; Tsybakov, 2009) given that the target function is Lipschitz continuous. Moreover, our theoretical findings align with convergence rates of nonparametric regression using deep neural networks (Bauer and Kohler, 2019; Nakada and Imaizumi, 2020; Schmidt-Hieber, 2020; Kohler and Langer, 2021; Farrell et al., 2021; Kohler et al., 2022; Jiao et al., 2023a). It is noteworthy that our results improve upon the rate ${\mathcal{O}}(n^{-\frac{2}{d+5}})$ derived by Chen et al. (2023b).
In Theorem 3.10, the hypothesis class $\mathscr{B}$} is set as a deep neural network class with Lipschitz constraints. The assumption of uniformly Lipschitz continuity of velocity estimator plays a crucial role in controlling the discretization error induced by Euler method. For further details, refer to Theorem 3.12. This assumption is standard in the theoretical analysis for flow-based or diffusion models, as considered by Kwon et al. (2022, Assumption A2) and Chen et al. (2023c, Assumption 3). In practical applications, various techniques for restricting the Lipschitz constant of deep neural networks have been proposed, such as weight clipping (Arjovsky et al., 2017), gradient penalty (Gulrajani et al., 2017), spectral normalization (Miyato et al., 2018), and Lipschitz network (Zhang et al., 2022). In the theoretical perspective, the approximation properties of deep neural network with Lipschitz constraint has been studied by Chen et al. (2022); Huang et al. (2022); Jiao et al. (2023b); Ding et al. (2024). In this work, an approximation error bound for deep neural networks with Lipschitz constraint is established in Section J.
@@ 3.4. Analysis for Euler Sampling
Under the same conditions as Theorem 3.10. Let the number of time stepss for Euler method $K$ be a positive integer. Then the following inequality holds
@@
The error bound of the Euler flow (2.9), as derived in Theorem 3.12, can be roughly divided into two main components. The first term arises from velocity matching, aligning with the error bound presented in Corollary 3.11. The second term corresponds to the discretization error introduced by Euler method. Moreover, as the number of time stepss $K$ increases, the error bound in Theorem 3.12 converges to that in Corollary 3.11.
@@
Under the same conditions as Theorem 3.12. Suppose the number of time stepss for Euler method $K$ satisfies $K\geq Cn^{\frac{1}{d+3}}$. Then the following inequality holds
@@
Corollary 3.13 presents a bound for the total error of flow-based generative models. The first term in the error bound corresponds to the 2-Wasserstein error of the Euler flow, as derived in Theorem 3.12. The second term captures the convergence of the interpolant distribution $\mu_{T}$ to the target distribution $\mu_{1}$. It is worth noting that as the stopping time $T$ approaches one, the first term tends to infinity, while the second term simultaneously decreases. This trade-off within the error bound highlights the importance of carefully selecting the stopping time $T$ and provides practical guidance for determining it in real-world applications. Since the optimal early-stopping time depends on the specific choice of the interpolant coefficients, $\alpha(t)$ and $\beta(t)$, we derive the optimal early-stopping time $T$ for two important cases: linear interpolant and Föllmer flow in Corollaries 3.16 and 3.17.
@@ 3.5. Analysis for Characteristic Generator
${\mathscr{G}}=\left\{g\in N(L,S):\begin{aligned} &\|g(t,s,x)\|_{\infty}\leq B_{\operatorname{flow}}\log^{1/2}m,\\ &\|\partial_{t}g(t,s,x)\|_{2},\|\partial_{s}g(t,s,x)\|_{2}\leq 3B^{\prime}_{\operatorname{flow}}\log^{1/2}m,\\ &\|\nabla g(t,s,x)\|_{\operatorname{op}}\leq 3\exp(GT),\penalty 10000\ 0\leq t\leq s\leq T,\penalty 10000\ x\in{\mathbb{R}}^{d}\end{aligned}\right\},$
where the depth and the number of non-zero weights of the neural network are given, respectively, as $L=C$ and $S=Cm^{\frac{d+2}{d+4}}$. Then it follows that
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{Z}}}\big[{\mathcal{D}}(\widehat{g})\big]\leq C\kappa^{2}(T)\Big\{n^{-\frac{2}{d+3}}\log^{2}n+\frac{\log n}{K^{2}}\Big\}+C\Big\{m^{-\frac{2}{d+4}}\log^{2}m+\frac{\log m}{K}\Big\},$
@@
(3.4) $K\geq\max\big\{Cn^{\frac{1}{d+3}},C\kappa^{-2}(T)n^{\frac{2}{d+3}}\big\}\quad\text{and}\quad m\geq C\kappa^{-(d+4)}(T)n^{\frac{d+4}{d+3}},$
@@
Remark 3.15.
The error bound in Theorem 3.14 is an average-in-time guarantee, not a uniform one. This is a direct consequence of our least-squares objective (2.5), which is designed to minimize the error averaged over all time pairs $(t,s)$. While this formulation is effective for learning the overall map, it naturally leads to a convergence guarantee in an averaged sense. However, it does not imply a stronger, uniform guarantee for a specific time pair. Some potential approaches to obtain a uniform convergence guarantee will be discussed in Section 6.
@@ 3.6. Applications
In this section, the theoretical analysis is applied to two types of flow-based method in Table 3: linear interpolant and Föllmer flow. The convergence rates of them are shown in Corollaries 3.16 and 3.17, respectively.
@@
Suppose Assumptions 1 and 2 hold. Set the velocity matching network class $\mathscr{B}$} to be Lipschitz-controlled, with depth $L_{b}=C$ and number of non-zero weights $S_{b}=Cn^{\frac{d+1}{d+3}}$. Set the stopping time $T$ as
@@
In addition, set the characteristic network class to be Lipschitz-controlled, with depth $L_{g}=C$ and number of non-zero weights $S_{g}=C\,m^{\frac{d+2}{d+4}}$. Further, let
$m\ \geq\ C\,n^{\frac{d+4}{3(d+3)}}\,\log^{\,d+4}\!n.$
Then it follows that
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{Z}}}\big[{\mathcal{D}}(\widehat{g})\big]\leq C\,n^{-\frac{2}{3(d+3)}}\,\log n.$
Here, the constants $C$ may depend on $d$ and $\sigma$, but is independent of $n$.
@@
Suppose Assumptions 1 and 2 hold. Set the velocity matching network class $\mathscr{B}$} to be Lipschitz-controlled, with depth $L_{b}=C$ and number of non-zero weights $S_{b}=Cn^{\frac{d+1}{d+3}}$. Set the stopping time $T$ as
@@
${\mathbb{E}}_{{\mathcal{S}}}\Big[W_{2}^{2}\Big((\widehat{E}_{0,K}^{\tau})_{\sharp}\mu_{0},\mu_{1}\Big)\Big]\leq C\,n^{-\frac{2}{5(d+3)}}\,\log^{\frac{3}{5}}n.$
In addition, set the characteristic network class to be Lipschitz-controlled, with depth $L_{g}=C$ and number of non-zero weights $S_{g}=C\,m^{\frac{d+2}{d+4}}$. Further, let
$m\ \geq\ C\,n^{\frac{d+4}{5(d+3)}}\,\log^{\,\frac{6}{5}(d+4)}\!n.$
Then it follows that
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{Z}}}\big[{\mathcal{D}}(\widehat{g})\big]\leq C\,n^{-\frac{2}{5(d+3)}}\,\log^{\frac{3}{5}}n.$
Here, the constants $C$ may depend on $d$ and $\sigma$, but is independent of $n$.
3.7. Mitigate the Curse of Dimensionality
As shown in Sections 3.10 and 3.14, the convergence of both velocity matching and characteristic fitting suffer from the curse of dimensionality: the sample complexity increases exponentially with the ambient dimension $d$ of the data. However, we will show that under Assumption 3, the convergence rate of the characteristic generator overcomes this limitation. The key to this result is the following decomposition lemma.
Lemma 3.18 (Low-dimensional decomposition).
Suppose Assumptions 1 and 3 hold. For each $(t,s,x)$ with $0\leq t\leq s\leq 1$ and $x\in{\mathbb{R}}^{d}$,
$b^{*}(t,x)$ $\equiv P\,\widetilde{b}^{*}\big(t,P^{\top}x\big)+\frac{\alpha_{t}\dot{\alpha}_{t}+\sigma^{2}\beta_{t}\dot{\beta}_{t}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\,(I_{d}-PP^{\top})x,$
$g^{*}(t,s,x)$ $\equiv P\,\widetilde{g}^{*}\big(t,s,P^{\top}x\big)+\sqrt{\frac{\alpha_{s}^{2}+\sigma^{2}\beta_{s}^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}}\,(I_{d}-PP^{\top})x.$
The vector field $\widetilde{b}^{*}:{\mathbb{R}}\times{\mathbb{R}}^{d^{*}}\rightarrow{\mathbb{R}}^{d^{*}}$ is defined as:
$\widetilde{b}^{*}(t,\widetilde{x}):={\mathbb{E}}[\dot{\alpha}_{t}\widetilde{X}_{0}+\dot{\beta}_{t}\widetilde{X}_{1}|\widetilde{X}_{t}=\widetilde{x}],\quad\widetilde{x}\in\mathbb{R}^{d^{*}},$
where $\widetilde{X}_{0}\sim N(0,I_{d^{*}})$ and $\widetilde{X}_{1}\sim N(0,\sigma^{2}I_{d^{*}})*\tilde{\nu}$ are independent, and $\widetilde{X}_{t}:=\alpha_{t}\widetilde{X}_{0}+\beta_{t}\widetilde{X}_{1}$. Further, the vector field $\widetilde{g}^{*}:{\mathbb{R}}\times{\mathbb{R}}\times{\mathbb{R}}^{d^{*}}\rightarrow{\mathbb{R}}^{d^{*}}$ is defined as the flow map of $\mathrm{d}\widetilde{x}(t)=\widetilde{b}^{*}(t,\widetilde{x}(t))\,\mathrm{d}t$.
This lemma is the cornerstone of our result. It reveals that both $b^{*}$ and $g^{*}$ have a crucial low-dimensional structure. They decompose into two orthogonal parts:
  1. (i)
    A complex, non-linear component governed by the low-dimensional functions $\widetilde{b}^{*}$ and $\widetilde{g}^{*}$, which only act on the $d^{*}$-dimensional data manifold.
  2. (ii)
    A simple, linear component that acts on the space orthogonal to the manifold.
This decomposition implies that the learning problem is greatly simplified. Instead of learning a function in the ambient $d$ dimensions, we only need to learn the low-dimensional components $\widetilde{b}^{*}$ and $\widetilde{g}^{*}$ in the intrinsic $d$ dimensions. This is precisely why our method’s convergence rate depends on $d^{*}$ instead of $d$, effectively mitigating the curse of dimensionality. The necessary regularity conditions for these low-dimensional functions are established in the appendix.
Now, we formally state the convergence results of linear interpolant and Föllmer flow under manifold hypothesis, which explicitly ease the curse of dimensionality.
Corollary 3.19 (Convergence of linear interpolant under manifold hypothesis).
Suppose Assumptions 1 and 3 hold. Set the velocity matching network class $\mathscr{B}$} to be Lipschitz-controlled, with depth $L_{b}=C$ and number of non-zero weights $S_{b}=Cn^{\frac{d^{*}+1}{d^{*}+3}}$. Set the stopping time $T$ as
$T=1-C\,n^{-\frac{1}{3(d^{*}+3)}}\,\log^{\frac{1}{2}}n.$
@@ 3.7. Mitigate the Curse of Dimensionality
Suppose the number of time step $K$ satisfies $K\geq C\,n^{\frac{1}{d^{*}+3}}$. Then it follows that
${\mathbb{E}}_{{\mathcal{S}}}\Big[W_{2}^{2}\Big((\widehat{E}_{0,K}^{\tau})_{\sharp}\mu_{0},\mu_{1}\Big)\Big]\leq C\,n^{-\frac{2}{3(d^{*}+3)}}\,\log n.$
In addition, set the characteristic network class to be Lipschitz-controlled, with depth $L_{g}=C$ and number of non-zero weights $S_{g}=C\,m^{\frac{d^{*}+2}{d^{*}+4}}$. Further, let
$m\ \geq\ C\,n^{\frac{d^{*}+4}{3(d^{*}+3)}}\,\log^{\,d^{*}+4}\!n.$
Then it follows that
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{Z}}}\big[{\mathcal{D}}(\widehat{g})\big]\leq C\,n^{-\frac{2}{3(d^{*}+3)}}\,\log n.$
Here, the constants $C$ may depend on $d$, $d^{*}$, and $\sigma$, but is independent of $n$.
Corollary 3.20 (Convergence of Föllmer flow under manifold hypothesis).
Suppose Assumptions 1 and 3 hold. Set the velocity matching network class $\mathscr{B}$} to be Lipschitz-controlled, with depth $L_{b}=C$ and number of non-zero weights $S_{b}=Cn^{\frac{d^{*}+1}{d^{*}+3}}$. Set the stopping time $T$ as
$T=1-C\,n^{-\frac{2}{5(d^{*}+3)}}\,\log^{\frac{3}{5}}n.$
Suppose the number of time step $K$ satisfies $K\geq C\,n^{\frac{1}{d^{*}+3}}$. Then it follows that
${\mathbb{E}}_{{\mathcal{S}}}\Big[W_{2}^{2}\Big((\widehat{E}_{0,K}^{\tau})_{\sharp}\mu_{0},\mu_{1}\Big)\Big]\leq C\,n^{-\frac{2}{5(d^{*}+3)}}\,\log^{\frac{3}{5}}n.$
In addition, set the characteristic network class to be Lipschitz-controlled, with depth $L_{g}=C$ and number of non-zero weights $S_{g}=C\,m^{\frac{d^{*}+2}{d^{*}+4}}$. Further, let
$m\ \geq\ C\,n^{\frac{d^{*}+4}{5(d^{*}+3)}}\,\log^{\,\frac{6}{5}(d^{*}+4)}\!n.$
Then it follows that
${\mathbb{E}}_{{\mathcal{S}}}{\mathbb{E}}_{{\mathcal{Z}}}\big[{\mathcal{D}}(\widehat{g})\big]\leq C\,n^{-\frac{2}{5(d^{*}+3)}}\,\log^{\frac{3}{5}}n.$
Here, the constants $C$ may depend on $d$, $d^{*}$, and $\sigma$, but is independent of $n$.
In contrast to the results in Corollaries 3.16 and 3.17, the convergence rates under the manifold assumption depend on the intrinsic dimension $d^{*}$ rather than the ambient dimension $d$. This demonstrates that our characteristic generator can mitigate the curse of dimensionality when the data possesses an intrinsic low-dimensional structure.
@@ 4.1.3. Characteristic fitting
8:  $\widehat{{\mathcal{R}}}_{\operatorname{loc}}(D_{{\mathcal{S}},\phi})\leftarrow\|D_{{\mathcal{S}},\phi}(t,t,X_{t})-X_{0}\|_{2}^{2}$.
@@
11:  $\widehat{{\mathcal{R}}}_{\operatorname{glo}}(D_{{\mathcal{S}},\phi})\leftarrow\|g_{s,T}^{\operatorname{off}}\circ g_{u,s}\circ g_{t,u}^{\operatorname{int}}(X_{t})-g_{s,T}^{\operatorname{off}}\circ g_{t,s}^{\operatorname{off}}(X_{t})\|_{2}^{2}$.
@@
Kim et al. (2024) proposed a similar method, but our approach differs from the CTM in two significant aspects. First, the integral scheme $g^{\operatorname{int}}$ employed by CTM is Euler-based. While Euler method coincides with the first-order exponential integrator for VE-ODE (Song et al., 2021c), its numerical stability cannot be guaranteed for general probability flow ODEs. In contrast, the exponential integrator used in our method may ease potential training instability as it fully exploits the semi-linearity of the ODE system. Secondly, CTM relies on GAN training in their approach, borrowing a pre-trained discriminator and treating $g$ as the generator. This reliance on GAN training may cause potential training instability and require extra training of the discriminator. Our method, however, does not require this additional GAN training burden. Detailed comparisons between characteristic generator and CTM can be found in Appendix N.